# Tensor Networks for Explainable Machine Learning in Cybersecurity

### 29 Dec 2023 v1

@@ II.1 Unsupervised generative modeling with MPS
Distinct from traditional generative models, this MPS framework can compute the partition function with linear complexity in the system size and boasts a direct sampling method that constructs samples incrementally. The model allows for deep theoretical understanding of its expressive power, adaptive adjustability, efficient computation of exact gradients and log-likelihood, and efficient direct sampling—position, making it therefore a powerful, efficient and interpretable tool for unsupervised learning and generative tasks.
@@ III Adversary-Generated threat intelligence
Our attack-detection system will identify attacks as anonalies. Notably, not all detected anomalies will be attacks. Indeed, there may be low-probability events that are perfectly fine (say, energy outages, reboots of the computer systems…). Because of this, it is important to play with the different parameters available in the model to minimize the detection of false positives, i.e., anomalies that are not attacks after all. With this in mind, central to the model’s success is the implementation of an acceptance threshold for the Negative Log Likelihood (NLL) of each event being generated by the model, so that events that exceed the threshold are classified as anomalous and may be considered by a human analyst to check if they correspond to a real attack. Establishing this threshold is challenging and varies in practice; some analysts may use statistical metrics like the mean and standard deviation to discern attacks within anomalies, while others may opt for the median. Through rigorous training, which leverages 70% of the data and respects the chronological order of events, our model effectively captures attack patterns. This temporal fidelity is crucial as shuffling the order of events has been shown to influence model training negatively. Notably, the average NLL of daily benign events is consistently lower than that of attacks, confirming the model’s adeptness at flagging attacks as anomalies.
@@ IV.2 Von Neumann entropy
being $\rho_{A}$ the RDM of subsystem $A$. In our setting, the entropy of a single MPS site measures the degree of entanglement between teh feature corresponding to that site andf the rest of the features in the system. In machine learning jargon, this is a measure of the contextual interdependencies of the features in the model.
@@ V Conclusions and Next Steps
In this paper we have discussed in detail the potential of TN methods to develop explainable and interpretable machine learning algorithms. We have shown the validity of the approach for an MPS algorithm designed for clustering, with the application to a real-life use case in cybersecurity. Our research has unveiled the potential of MPS as a powerful tool for unsupervised anomaly detection, particularly within the domain of adversary-generated threat intelligence. The key findings demonstrate that MPS not only rivals but in certain respects surpasses the performance of traditional machine learning models such as autoencoders and GANs, especially in terms of explainability and interpretability. The application of MPS to a cybersecurity dataset revealed its robustness in identifying cyberthreats with a low false positive rate, a critical parameter in cybersecurity operations. The ability of MPS to provide direct probability extraction, assess feature importance, and elucidate on the reasons behind the classification of data points as anomalies, underscores its utility in operational environments where understanding the ‘why’ behind a model’s decision is as vital as the decision itself. In addition, the research has highlighted the interpretability capabilities of MPS, offering insights into Von Neumann entropy, feature importance, conditional probabilities, and mutual information directly from the model’s parameters without the need for additional tools or techniques. These capabilities have profound implications for explainable AI, setting a new benchmark for transparency in machine learning applications.
The methods developed in this work can be improved in many ways. For instane, we could consider more advanced TN structures beyond MPS, with more intrincate patterns of correlations. We leave this for future studies.

### 05 Apr 2024 v3

@@ II.1 Unsupervised generative modeling with MPS
Distinct from traditional generative models, this MPS framework can compute the partition function with linear complexity in the system size and boasts a direct sampling method that constructs samples incrementally. The model allows for deep theoretical understanding of its expressive power, adaptive adjustability, efficient computation of exact gradients and log-likelihood, and efficient direct sampling-position, making it therefore a powerful, efficient and interpretable tool for unsupervised learning and generative tasks.
@@ III Adversary-Generated threat intelligence
Our attack-detection system will identify attacks as anomalies. Notably, not all detected anomalies will be attacks. Indeed, there may be low-probability events that are perfectly fine (say, energy outages, reboots of the computer systems…). Because of this, it is important to play with the different parameters available in the model to minimize the detection of false positives, i.e., anomalies that are not attacks after all. With this in mind, central to the model’s success is the implementation of an acceptance threshold for the Negative Log Likelihood (NLL) of each event being generated by the model, so that events that exceed the threshold are classified as anomalous and may be considered by a human analyst to check if they correspond to a real attack. Establishing this threshold is challenging and varies in practice; some analysts may use statistical metrics like the mean and standard deviation to discern attacks within anomalies, while others may opt for the median. Through rigorous training, which leverages 70% of the data and respects the chronological order of events, our model effectively captures attack patterns. This temporal fidelity is crucial as shuffling the order of events has been shown to influence model training negatively. Notably, the average NLL of daily benign events is consistently lower than that of attacks, confirming the model’s adeptness at flagging attacks as anomalies.
@@ IV.2 Von Neumann entropy
being $\rho_{A}$ the RDM of subsystem $A$. In our setting, the entropy of a single MPS site measures the degree of entanglement between the feature corresponding to that site and the rest of the features in the system. In machine learning jargon, this is a measure of the contextual interdependencies of the features in the model.
@@ V Conclusions and Next Steps
In this paper we have discussed in detail the potential of TN methods to develop explainable and interpretable machine learning algorithms. We have shown the validity of the approach for an MPS algorithm designed for clustering, with the application to a real-life use case in cybersecurity. Our research has unveiled the potential of MPS as a powerful tool for unsupervised anomaly detection, particularly within the domain of adversary-generated threat intelligence. The key findings demonstrate that MPS not only rivals but in certain respects surpasses the performance of traditional machine learning models such as autoencoders and GANs, especially in terms of explainability and interpretability. The application of MPS to a cybersecurity dataset revealed its robustness in identifying cyberthreats with a low false positive rate, a critical parameter in cybersecurity operations. The ability of MPS to provide direct probability extraction, assess feature importance, and elucidate on the reasons behind the classification of data points as anomalies, underscores its utility in operational environments where understanding the ‘why’ behind a model’s decision is as vital as the decision itself. In addition, the research has highlighted the interpretability capabilities of MPS, offering insights into Von Neumann entropy, feature importance, conditional probabilities, and mutual information directly from the model’s parameters without the need for additional tools or techniques. These capabilities have profound implications for explainable AI, setting a new benchmark for transparency in machine learning applications. In addition, the methods presented in this paper allow in the capability to generate synthetic data in order to improve training of complex models (due to lack of data availability) and activity generation for deception systems. This would help also in the development of behavior-based models, which are better against unknown attacks than rule-based models.
The methods developed in this work can be improved in many ways. For instace, we could consider more advanced TN structures beyond MPS, with more intrincate patterns of correlations. We leave this for future studies.
@@
Data availability statement: authors are open to reasonable requests.